OpenStudy (anonymous):

Let g be a function defined by g(x)=Integral from -1 to x of [(t^3-t^2-6t)/√t^2 +7] On which of the following intervals is g decreasing? A. x≤-2 and 0≤x≤3 B. x≤-2 and x≥3 C. -2≤x≤0 and x≥3 D. -2≤x≤3 E. x≤-1 I'm confused as to whether this function actually includes an integral or not.

jimthompson5910 (jim_thompson5910):

Just to be clear, the g(x) function is this right? \[\Large g(x) = \int_{-1}^{x}\frac{t^3-t^2-6t}{\sqrt{t^2+7}}dt\]

jimthompson5910 (jim_thompson5910):

If so, then you use the second fundamental theorem of calculus http://mathworld.wolfram.com/FundamentalTheoremsofCalculus.html to go from this \[\Large g(x) = \int_{-1}^{x}\frac{t^3-t^2-6t}{\sqrt{t^2+7}}dt\] to this \[\Large g \ '(x) = \frac{x^3-x^2-6x}{\sqrt{x^2+7}}\] you'll use `g ' (x)` to determine where g(x) is increasing or decreasing

OpenStudy (anonymous):

Huh

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

Un momento señor

OpenStudy (anonymous):

So, to determine if it's decreasing, I would just have to see if the slope is negative, right?

jimthompson5910 (jim_thompson5910):

which is equivalent to seeing where `g ' (x)` is negative

jimthompson5910 (jim_thompson5910):

The denominator of `g ' (x) ` is never negative. So you just need to focus on the numerator

OpenStudy (anonymous):

So THAT's the issue. I was looking at the function's slope, but not as a derivative function itself...

jimthompson5910 (jim_thompson5910):

The derivative represents the slope of the tangent lines, so they are related ideas.

OpenStudy (anonymous):

So it'd be A

jimthompson5910 (jim_thompson5910):

I agree. The answer is A

OpenStudy (anonymous):

Thanks mate

jimthompson5910 (jim_thompson5910):

no problem